OpenStudy (anonymous):

how to simplify this division radical 3 squared 20 / 2 ???

OpenStudy (anonymous):

Do you mean \[3\sqrt{20}/2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay this is simple, we must look at the 20 under the square root and say that it can be written as 4 time 5, then: \[3\sqrt{4*5}/2\] Next, the 4 comes out of the radical as a 2; \[3*2\sqrt{5}/2\] Then, we have: \[6\sqrt{5}/2\] Finally, the we can simplify this as; \[3\sqrt{5}\]

OpenStudy (anonymous):

okay thank you :)